OpenStudy (anonymous):

A solution of H2SO4(aq) with a molal concentration of 4.80 m has a density of 1.249 g/mL. What is the molar concentration of this solution?

OpenStudy (anonymous):

suppose mass of solvent 1 kg...(1000g) then molality =n(solute) / m(solvent) => 4.8m = 4.8mole / 1kg m(H2SO4) = n(H2SO4)*Mr(H2SO4) => m(H2SO4) = 4.8 mole * 98 g/mole = 470g m(solution) = m(H2SO4) + m(solvent) = 470 + 1000 = 1470 g d(solution) = m(solution) / V(solution) => => 1.249 g/mL = 1470 g / V(solution) => => V(solution) = 1470 g / 1.249 g/mL = 1177 mL ~ 1.2 L Molarity = n(solute) / V(solution) => => Molarity = 4.8 mole / 1.2 L = 4 M (there may be some mistakes..)